# Coach not crutch: AI assistance can enhance rather than hinder skill development

### v3

@@ Appendix C Results Study 2
To test for moderation effects of pre-treatment demographic variables, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that practicing with AI was equally effective across groups.
@@ Appendix D Results Study 3
D1 Mediation
Appendix E Results Study 4
E1 Randomization, Balance, and Missingness
Participants in Study 4 were recontacted to return two days after the initial intake section. While most people responded, 14% of recontacted participants did not respond. As shown in Table S15, missingness and attrition rates were low and did not differ by condition.
@@ Appendix E Results Study 4
E2 Details on professional editors
@@
To determine if editors changed their behavior over time, we fit a series of multilevel models, with cover letter position as a predictor, random intercept for editors, and text quality (as rated by GPT-4o), time spent per cover letter, and number of insertions, deletions, or modifications (i.e., Levenstein distance). These analyses include editors who were able to edit at least 20 letters, and focus on the first 20 letters edited. The quality of the edited letters improved over time ( = 0.013, SE = 0.005, p = 0.013), even though editors spent less time as the task progressed ( = -0.387, SE = 0.028, p < .001). Interestingly, editors also were able to improve the emails using fewer edits as the task progressed ( = -2.572, SE = 0.926, p = 0.00561). Despite spending less time per letter, editors produced higher quality edits with fewer but more effective changes, demonstrating clear learning effects in professional editing tasks.
E3 Details on participants’ search behavior
@@
Across 635 participants assigned to use Google to search for cover letter examples and tips, we found search behavior consistent with prior research on web-browsing behavior (39). Most participants (97%) stayed on page 1, while only 3% ventured beyond the first page of results. Participants spent an average of 2.8 minutes (median: 1.3 minutes) browsing search results, with an average of 0.9 minutes between clicks. On average, participants made 2.9 clicks during their search session.
E4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
@@
E5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
@@
E6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
@@
E7 Results were not moderated by individual differences
As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example. To do this, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups. See Table S22
@@
Appendix F Results Study 5
F1 Randomization, Balance, and Missingness
@@ Appendix F Results Study 5
F2 AI examples improve writing skill
@@
F3 Seeing AI examples was less effortful
@@
Table S29 shows OLS models predicting test effort metrics from practice condition. Results show some differences: participants who had seen the AI example write for longer duing the test, and pressed more keys, however their subjective experience of effort was not different from those who practice with or without the AI tool.
@@
F4 Seeing an AI example did not discourage motivation for future learning
@@
F5 The benefits of seeing an AI example were just as large a day later
@@
F6 Seeing AI examples was equally effective across subgroups
As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example. To do this, we ran separate linear in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups.

### v4

@@ Appendix C Results Study 2
To test for moderation effects of pre-treatment demographic variables, we ran separate linear regressions in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that practicing with AI was equally effective across groups.
@@
C8 Forecasting
| Predictor | OR | SE | z | p-value |
| Prior AI experience | 0.718 | 0.199 | -1.664 | 0.096 |
| Age (years) | 1.014 | 0.021 | 0.675 | 0.500 |
| Male (vs. Female) | 1.035 | 0.467 | 0.073 | 0.942 |
| White (vs. Other) | 0.966 | 0.509 | -0.068 | 0.946 |
| Student (vs. Not) | 1.120 | 0.618 | 0.183 | 0.855 |
| Employment status | ||||
| Other | 0.770 | 0.687 | -0.382 | 0.703 |
| Part-time | 0.663 | 0.601 | -0.684 | 0.494 |
| Unemployed (seeking) | 0.645 | 0.593 | -0.738 | 0.461 |
@@ Appendix D Results Study 3
D1 Randomization, Balance, and Missingness
Participants in Study 3 were recontacted to return two days after the initial intake section. While most people responded, 14% of recontacted participants did not respond. As shown in Table S17, missingness and attrition rates were low and did not differ by condition.
@@
D2 Details on professional editors
@@
To determine if editors changed their behavior over time, we fit a series of multilevel models, with cover letter position as a predictor, random intercept for editors, and text quality (as rated by GPT-4o), time spent per cover letter, and number of insertions, deletions, or modifications (i.e., Levenshtein distance). These analyses include editors who were able to edit at least 20 letters, and focus on the first 20 letters edited. The quality of the edited letters improved over time ( = 0.013, SE = 0.005, p = 0.013), even though editors spent less time as the task progressed ( = -0.387, SE = 0.028, p < .001). Interestingly, editors also were able to improve the emails using fewer edits as the task progressed ( = -2.572, SE = 0.926, p = 0.00561). Despite spending less time per letter, editors produced higher quality edits with fewer but more effective changes, demonstrating clear learning effects in professional editing tasks.
Why did participants learn more from AI than they did from professional human editors. One possibility is that the quality of the edits provided by AI were of higher quality than those of human editors. To test this possibility, we used our AI tool to rewrite the same letters that were assigned to professional human editors. That way, for each participant in the “practice with editor feedback condition” we have the actual letter they received (as edited by the human expert), and the counterfactual letter they would have received had they been assigned to practice with AI. As shown in Figure 8(a), a sample of naive human evaluators were twice as likely to prefer hte cover letter edited by AI than the one participants received, edited by a professional human editor (paired d = 0.62, p < .001) Figure 8(b) shows the densities of writing quality as evaluated by our AI rating procedure. Cover letters edited by AI were rated as significantly higher quality than those submitted by professional human editors (paired d = 1.66, p < .001)
D3 Details on participants’ search behavior
@@
Across 635 participants assigned to use Google to search for cover letter examples and tips, we observed search behavior consistent with prior research on web browsing (pan2007google). Most participants (97%) stayed on page 1, while only 3% ventured beyond the first page of results. Participants spent an average of 2.8 minutes (median: 1.3 minutes) browsing search results, with an average of 0.9 minutes between clicks. On average, participants made 2.9 clicks during their search session.
D4 AI practice improved writing skill more than getting feedback from professional editors and looking for examples online
@@
D5 AI practice wasn’t any more effortful than getting feedback from professional editors and was less effortful than looking for examples online
@@
D6 AI practice did not create more of an illusion of mastery than getting feedback from professional editors or looking for examples online
@@
D7 Results were not moderated by individual differences
As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example. To do this, we ran separate linear regressions in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups. See Table S24
@@
D9 Forecasting and Willingness to Pay
Appendix E Results Study 4
E1 Randomization, Balance, and Missingness
@@ Appendix E Results Study 4
E2 AI examples improve writing skill
@@
E3 Seeing AI examples was less effortful
@@
Table S31 shows OLS models predicting test effort metrics from practice condition. Results show some differences: participants who had seen the AI example wrote for longer during the test, and pressed more keys; however, their subjective experience of effort was not different from those who practiced with or without the AI tool.
@@
E4 Seeing an AI example did not discourage motivation for future learning
@@
E5 The benefits of seeing an AI example were just as large a day later
@@
E6 Seeing AI examples was equally effective across subgroups
As in Study 2, we tested whether each of the pretreatment demographic variables moderated the effects of seeing an AI example. To do this, we ran separate linear regressions in which writing skill during the test phase was regressed on condition, the pre-treatment moderator of interest, writing skill at baseline, and an interaction term between the moderator condition. After correcting the p-values for the interaction terms, none were significant at the .05 level, suggesting that seeing AI examples was equally effective across groups.